OpenStudy (anonymous):

Find the slope of the line whose equation is 4x – 6y = 12.

OpenStudy (tkhunny):

Have you considered solving for y?

OpenStudy (anonymous):

ha i didn't seee that brb

OpenStudy (anonymous):

i got -2

OpenStudy (tkhunny):

4x – 6y = 12 Divide by 2 just to make life simpler. 2x – 3y = 6 Subtract 2x –3y = -2x + 6 Divide by -3 y = (2/3)x - 2 Solving for y produces the slope-intercept form!! There's the slope you need! (0,-2) is the y- intercept. The slope is the other number.

OpenStudy (anonymous):

thank you i bee stuck on this problem for 15 min's

OpenStudy (tkhunny):

Did you get 2/3?

OpenStudy (johnmistrider02):

Thanks